Ifcos x = 3 cos y  , then 2 tan (y/2-x/2) =
a).cot (y/2-x/2) b). cot(x/4+y/4) c). cot (y/4-x/4) d). cot(x/2+y/2)

Grade:11

1 Answers

Sankalp Das
6 years ago
cosx/cosy =3applying componendo and dividendo(Cosx+cosy)/(cosx-cosy)=3+1/3-1{2cos(x/2+y/2)cos(x/2-y/2)}/{-2sin(x/2+y/2)sin(x/2-y/2)=2Cot(x/2+y/2)=2(-tan(x/2-y/2))cot(x/2+y/2)=2tan(y/2-x/2)(d)is correct